If 100ft of atmospheric air at zero Fahrenheit temperature are compressed to a volume of 1ft at a temperature of 200F. what will be the pressure of the air in psi? Ans. 2109 psia (EE board problem)


1
Expert's answer
2021-11-29T11:46:21-0500

"\frac{p_1V_1}{T_1}=\frac{p_2V_2}{T_2}, \implies"

"p_2=\frac{p_1V_1T_2}{T_1V_2}=2019~\text{psi}."